OpenStudy (anonymous):

What is the equation of the ellipse with foci at (5, 0), (-5, 0) and vertices (9, 0), (-9, 0)? Answers: x2/81+y2/25=1 x2/81-y2/25=1 x2/81+y2/56=1 x2/81-y2/56=1

OpenStudy (anonymous):

@yrelhan4

OpenStudy (yrelhan4):

coordinates of focii are (+/-ae,0).. and coordinates of vertices are.. (=/-a,0).. and e=sqrt(1-[(b/a)]^2).. using the first two. find the value of e.. once you have e.. using the third relation find the value of b.. once you have a and b.. you can easily check which equation satisfies..

OpenStudy (yrelhan4):

where 2a=length of major axis.. 2b=length of minor axis.. e=eccentricity. i hope you know these terms..

OpenStudy (anonymous):

yea I do My answer that I got is A

OpenStudy (yrelhan4):

hmm. give me a minute.

OpenStudy (yrelhan4):

i am not able to solve it for some reason.. lets solve together and tell me where i'm going wrong..

OpenStudy (yrelhan4):

so.. ae=5 and a=9.. right?

OpenStudy (anonymous):

yea

OpenStudy (yrelhan4):

e=5/9=sqrt(1-(b^2/a^2))..

OpenStudy (yrelhan4):

--> b^2/a^2=1-25/81..

OpenStudy (yrelhan4):

=56/25 --> b^2=56..

OpenStudy (yrelhan4):

oh wait.. oh i did that wrong..

OpenStudy (yrelhan4):

b^2/a^2=56/81 -->b^2=56..

OpenStudy (yrelhan4):

so your answer should be C..

OpenStudy (anonymous):

I see where I made my mistake Thank You

OpenStudy (yrelhan4):

Cheers.